# A Bio-Inspired Research Paradigm of Collision Perception Neurons Enabling Neuro-Robotic Integration: The LGMD Case

### v1

@@ 3 LGMD-based dynamic vision systems facilitate collision-free navigation
Although these models are established with biologically plausible formulations, they are rarely used to address real-world challenges directly owing to their reliance on the physical attributes of looming objects that are impossible to measure in unpredictable, real-world environments. To alleviate this, Keil developed a multi-layer network that uses luminance values, captured by CCD or CMOS cameras, as model input. The response of this model exhibits characteristics similar to those of the $\eta$-function, without prior knowledge on image size or velocity [59]. This model was subsequently extended with ON/OFF pathways to enhance its performance across various real-world scenes [60]. However, the computing units in these models are formulated using ordinary differential equations, which makes the computations time-consuming and limits their applicability in robotics for real-time navigational applications. Additionally, these approaches employ hierarchical structures similar to the four-layer LGMD1 network proposed by Rind, rather than the single-neuron computation represented by the $\eta$-function.
@@
In contrast, the four-layer LGMD neural network proposed by Rind et al. appears particularly attractive to robotic engineers. The first implementation of this four-layer LGMD network for collision detection in ground mobile robots occurred in 2000, where a Khepera mobile robot equipped with a CCD camera was used to perceive real-world environments. Collision detection and avoidance were achieved using the simulation software IQR421 [10]. Subsequently, Yue et al. combined genetic algorithms with the LGMD neural network to address collision detection in car driving scenarios [62]. Further improvements were made by incorporating an artificial grouping layer into the four-layer LGMD network, thereby enhancing its performance when managing complex real-world tasks [61] (See Fig. 5B and Fig. 6A). Building on similar ideas, Yue and Rind continued exploring the potential of LGMD-based near-range path navigation, including the integration of panoramic images and regular CCD images for improved collision detection [63, 64]. Moreover, Silva et al. extended the LGMD model by integrating previous work from [61] and [65]. They verified a linear relationship between time-to-collision444time difference between peak firing time of LGMD model (predicted time) and genuine collision time and time-of-collision, then investigated the $l/|v|$ ratio in MATLAB simulations and in online DKRK8000 robot navigation [66].
Another milestone was established in 2017 when the LGMD-based neural network model was built for the first time into the embedded vision system. Hu et al. adapted a similar vision-based LGMD network as in [61] into a small autonomous mobile robot called “Colias” [67]. The robot performance was validated through autonomous navigation in an arena mixed with obstacles (see Fig. 6C), demonstrating the reliability and robustness of the embedded vision system of LGMD case. Later, Fu et al. further facilitated collision-free navigation for Colias robot by gradually incorporating spike frequency adaptation, ON and OFF visual pathways, as well as feedback neural computation [13, 68, 69] (See Fig. 5C). Zhao et al. also utilized LGMD-based vision detectors to achieve collision-free flight of a small quadcopter (see Fig. 6D)[70, 71]. More recent studies have demonstrated that the LGMD-based vision system also holds great promise for boosting the safe navigation of hexapod walking robots and UAVs mounted with dynamic vision sensor (see Fig. 6E, F) [72, 15, 73].
@@
Furthermore, engineers and computational modelers have explored ways to enhance the looming selectivity of LGMD1-based perception methods by incorporating various computational mechanisms. These methods can be broadly classified into two categories: the first approach involves integrating mechanisms such as ON/OFF competition or directional sub-networks to suppress responses to receding or translating stimuli [78, 40]. The second approach employs binocular vision or additional units to estimate changes in the depth of moving objects, thereby distinguishing between approaching, receding, and translational stimuli [65, 79, 80]. However, these mechanisms increase computational burden, not yet to be validated in real-world robotic scenarios.
On the other hand, due to limited biological research on LGMD2 and its selective response to only OFF contrast looming stimuli, LGMD2 has relatively very limited numerical modeling and applications in robotics. The seminal work on LGMD2 modeling appeared late in 2015 when Fu and Yue introduced ON and OFF visual pathways and non-linearity to simulate the LGMD2’s selectivity for OFF contrast stimuli [31].
@@
The LGMD2 neural network model was optimized and smoothly embedded into the Colias micro-robot, achieving a collision detection and avoidance success rate of 95.3% in arena tests [81]. As LGMD1 and LGMD2 have different selectivity that could benefit addressing variable real-physical collision challenges, the first hybrid LGMD1-LGMD2 neural model was proposed to implemented as embedded vision system for micro-robot navigation [82]. In this research, the robot’s LGMDs vision behaved similarly to biological neurons that demonstrated the effectiveness of LGMD1 and LGMD2 together for artificial collision-detecting system. Unlike LGMD1, LGMD2 possesses a single large dendritic field, leading to the assumption that FFI is absent in its pre-synaptic circuitry. To accurately simulate the LGMD2’s selectivity, Fu et al. introduced an adaptive inhibition mechanism as a substitute for FFI, thereby aligning with physiological findings and enhancing the performance of the LGMD2-based visual system in complex real-world tasks [32, 14, 83].
@@
Modelers and engineers are actively seeking opportunities to extend LGMD research to new application scenarios and sensor technologies. For example, LGMD-based collision detection methods have been combined with path integration mechanisms discovered in bees, enabling collision-free autonomous navigation toward specific goal locations, even in the presence of both static and moving obstacles [84, 85, 86]. Unlike conventional cameras, event-based cameras detect changes in light intensity asynchronously, producing sparse, high-temporal-resolution data with low power consumption and a wide dynamic range [87]. This makes event-based cameras ideal inputs for ON/OFF pathways. When interpreted with spiking neural networks (SNN), event camera-based LGMD models compensate for the limitations of traditional RGB cameras, effectively detecting ultra-fast looming objects or targets in complex and diverse low-light environments, even facilitating the control of UAVs during flights [88, 89, 77]. Unlike previous methods that relied on normalization mechanisms to enhance model responses to low-contrast stimuli, Zhang et al. recently introduced thermal cameras to improve the performance of LGMD-based models. By leveraging thermal imaging, the ability to detect approaching objects was significantly enhanced, providing an alternative solution to detect collisions in challenging lighting conditions [90].
@@
In summary, LGMD-based collision detection methods provide a distinctive, data-free, and visually driven approach, standing in stark contrast to deep learning and sensor-based techniques. While deep learning models achieve remarkable accuracy by identifying intrinsic patterns within large datasets, they demand significant computational resources, extensive data, additional sensors, and lengthy training processes. In comparison, LGMD-based methods leverage biologically plausible neural structures to enable parsimonious and efficient collision detection without requiring high-performance hardware or large datasets. This makes them particularly well-suited for real-time applications in small mobile robots, where timeliness, computational simplicity, and energy efficiency are paramount. Despite these advantages, LGMD-based models face notable challenges in robustness when applied to diverse real-world scenarios. They lack the adaptability and flexibility of real LGMD neurons, especially in handling the complexities of dynamic and noisy environments. Addressing these limitations presents significant opportunities for future research to further refine and expand LGMD-based visual systems.
@@ 4 How these methods feedback to neuroscience
The behaviors of LGMD-based robotics in complex real-world interaction tasks have directly validated the robustness of the looming selectivity as artificial life comparable to that observed in LGMD neurons. In particular, the computational model of LGMD2 has been integrated into the micro-robot - Colias, demonstrating a selective looming response to darker objects during real-world navigation [32]. Additionally, the Strider robot, implemented with a four-layer LGMD1 network incorporating a different correlation framework, revealed a linear relationship between time to collision and the size-to-speed ratio ($l/|v|$), consistent with previous findings by Gabbiani et al. [36, 37]. Similarly, the Khepera robot, controlled by an LGMD-based collision detection and avoidance system, exhibited comparable behavior—specifically, as the moving speed increased, the robot generated collision warnings earlier, thereby maintaining a greater distance from obstacles [61]. F. Claire Rind, a neuroscientist who has dedicated several decades to studying the LGMD neurons, remarked that “Insects, and locust looming detectors in particular, have already provided inspiration for visual control of robots, unmanned autonomous vehicles, and aerial drones.” in a recent discussion on insects’ 3D vision [33]. Her insights underscore the profound impact of bio-inspired research in bridging the gap between neuroscience and technological innovation, demonstrating how understanding simple biological systems can drive advancements in robotics and autonomous systems.
@@
Bio-inspired robotics have the potential to contribute far beyond mere “verification”. Two decades ago, scholars proposed that bio-robotic studies could serve as effective paradigms for understanding animal behavior, suggesting that the performance of bio-inspired robotics might also offer valuable insights into neuroscience and behavior [91]. Regarding LGMD2, the strongest feedback to neuroscience is the existence of ON/OFF-contrast encoding neurons or neural pathways prior to the dendrites of LGMDs. Unlike research into fly visual systems, there is very limited evidence for whether/where the ON/OFF channels exist in locust’s visual circuitry. Considering the homology between different insects’ visual brains, and the universal findings of polarity vision across researched animals’ dynamic vision systems, the simulation of LGMD2’s circuits provides strong hypothesis back to neuroscience [32, 24]. At this point, Rind also commented “Simulations show how response tuning for dark transitions in the LGMD2 could complement LGMD1 responses [32].”
@@
In recent studies, the competition between ON/OFF channels also indicate how opposite-polarity signals could interact to suppress translating-induced excitation in order to sharpen up the looming selectivity of LGMD [40]. Furthermore, the performance of LGMD-based real-world applications is partially restricted in noisy and low-contrast environments. Keil’s computational model suggested that threshold-involved neuronal signal integration could pool signals from different channels, thereby enhancing looming selectivity in noisy environments [58]. By analyzing realistic synaptic integration within bio-physically accurate neuronal models, Alon found that introducing an additional threshold for dendritic spikes significantly increased neuronal tolerance to a wide variety of noise [92]. Besides, the recent studies proposed that normalizing photoreceptor signals can enhance LGMD responses to low-contrast stimuli [83, 93]. Similarly, global inhibition observed in the locust’s visual pathways helps maintain excitation within the dynamic range of the neuron, thereby preserving functional selectivity under varying illumination conditions [49]. These mechanisms are also present in other sensory circuits, highlighting their fundamental role in neural processing [94]. Moreover, the deepened LGMD1-LGMD2 cascade network shows selectivity to only approaching targets which might reveal some cooperative mechanisms interacting the pre-synaptic dendrites between LGMD1 and LGMD2, nevertheless unknown in terms of physiology [83, 11].
@@ 5 Discussion
Application demands - First, there is an urgent need for efficient, low-energy collision detection methods in real-world applications. Collisions are ubiquitous and pose significant hazards to both living creatures and intelligent robotic systems. Current collision detection methods are often energy-intensive due to the need for map reconstruction and object recognition [95, 96, 97], or they rely heavily on specific sensors such as radar or wireless ultraviolet light [98, 99]. Insects, with their tiny brains and highly effective collision avoidance behaviors, offer an ideal model for developing parsimonious and efficient collision detection systems. The evolved, specialized structure of looming-selective neural circuits in insects provides a foundation for designing straightforward models capable of addressing real-world collision challenges. There are two identified types of neurons that respond selectively to approaching objects: the LGMDs and the lobula plate/lobula columnar type-2 (LPLC2) neurons. LPLC2 neurons are located in the lobula complex of Drosophila, and unlike LGMD, their excitation relies on the precise spatial arrangement of dendrites within a directionally selective motion feature map provided by T4/T5 neurons [100, 101]. LPLC2 neurons prefer motion that is outward-from versus inward-toward their receptive field center, and this radial motion opponency causes LPLC2 to be excited by looming objects at the center of the receptive field rather than by receding objects. However, this characteristic limits LPLC2’s response in near-miss scenarios or to looming patterns from different directions, making it potentially insufficient for handling complex real-world challenges. In contrast, LGMD neurons can detect imminent collisions from various directions in both stationary and moving backgrounds, without requiring directionally specific afferent signals [30]. The robustness of LGMD responses makes it a suitable alternative for generating accurate yet efficient collision detection methods.
@@
Biological prominence - Second, comprehensive research in neuroscience and anatomy on the LGMD neuron has significantly contributed to the successful development of artificial systems. The LGMD neuron is large in size, with an extensive dendritic structure that makes it easier to study using various experimental protocols. Dendritic field A spans approximately $7705\mu m$ in length and $108,400\mu m^{2}$ in surface area, while dendritic fields B and C are shorter, with lengths of approximately $4438\mu m$ and $2517\mu m$, surface areas of $28,100\mu m^{2}$ and $13,500\mu m^{2}$, and radii of $0.8\mu m$ and $0.7\mu m$, respectively. This morphology facilitates research using tools such as extracellular hook electrodes, two-photon microscopy, and optogenetics [18, 102, 103, 104]. Furthermore, investigating the neuronal response of the LGMD is particularly feasible in laboratory settings because approaching objects can be effectively simulated using two-dimensional projections. These factors have sustained the vibrant wave of neuroscience and anatomical research on LGMD neurons since their discovery in the 1970s. Over time, this research has become increasingly comprehensive, addressing aspects ranging from morphology to membrane conductance [105, 106]. Consequently, it has provided an abundant source of inspiration for developing LGMD-based applications in real-world scenarios.
@@
Although the neuronal response of LGMDs can be effectively simulated using the NEURON simulation environment [^5] https://www.neuron.yale.edu/neuron/ [20], the afferent mechanisms driving LGMD responses to different types of motion remain largely unexplored. Olson et al. focused on replicating the neural signal processing of LGMD by incorporating global, lateral, and feedforward inhibition, demonstrating alignment between model outputs and real LMC responses. However, their model was limited in scope, considering only the OFF pathway while excluding the ON pathway, ON/OFF interactions, as well as self-inhibition and lateral excitation [22]. This incomplete representation highlights the gaps that persist in our understanding of the full processing dynamics of LGMDs. Moreover, the connectivity between LGMD neurons and other neurons within the locust neuropile remains largely undefined. Recent connectomic data from Drosophila suggested that integrating whole-field looming detection pathways with local directional signals can significantly enhance collision detection accuracy in both simulated and real-world environments [117, 118]. Additionally, modelers have combined LGMD1 with LPLC2 to improve looming perception performance in real-world scenarios, showing significant promise for high-speed collision detection tasks [119, 120]. Another unresolved question concerns the evolutionary rationale behind the emergence of distinct LGMD1 and LGMD2 neurons, each with unique morphology. Understanding why and how these specialized neurons developed over thousands of years is critical for replicating their adaptability and efficiency in engineered models.
@@
While we are still far from fully understanding the intrinsic mechanisms that enable the locust cerebellum to achieve its remarkable collision detection capabilities, even a simple replication of the LGMD neuronal response characteristics has already proven sufficient for developing low-energy, robust collision detection methods. This bio-inspired paradigm exemplifies a fundamentally different methodology compared to deep learning-based collision detection strategies. Deep learning methods, though originally inspired by brain functions, primarily emphasize the learning process. They derive their strength from uncovering intrinsic patterns and features within extensive training datasets, achieving outstanding performance through experience-based adaptation. In contrast, LGMD-based collision detection methods focus on directly replicating neuronal responses and neural architectures, providing an alternative approach for real-world collision detection that operates without relying on learning or large datasets. This bio-inspired paradigm of LGMD offers a promising pathway for efficient collision detection, especially in scenarios where simplicity, low computational overhead, and rapid responses are crucial. We believe these two paradigms—deep learning-based and LGMD-based approaches—can complement and enhance each other. Deep learning methods can provide information about approaching objects and the agent’s self-location, further enhancing the performance of LGMD-based collision detection methods, particularly in complex real-world scenarios. Conversely, LGMD-based methods can act as supplementary mechanisms to support deep learning models in situations where training data are limited or unavailable. Together, these approaches form a synergistic framework for advancing collision detection technologies.

### v2

@@ 3 LGMD-based dynamic vision systems facilitate collision-free navigation
3.1 LGMD1-based collision detectors for ground and aerial robots
@@
Although these models are established with biologically plausible formulations, they are rarely used to address real-world challenges directly owing to their reliance on the physical attributes of looming objects that are impossible to measure in unpredictable, real-world environments. To alleviate this, Keil developed a multi-layer network that uses luminance values, captured by CCD or CMOS cameras, as model input. The response of this model exhibits characteristics similar to those of the $\eta$-function, without prior knowledge on image size or velocity [60]. This model was subsequently extended with ON/OFF pathways to enhance its performance across real-world scenes [61]. However, the computing units in these models are formulated using ordinary differential equations, which makes the computations time-consuming and limits their applicability in robotics for real-time visual processing and navigational applications. Additionally, these approaches employ hierarchical structures similar to the four-layer LGMD1 network proposed by Rind, rather than the single-neuron computation represented by the $\eta$-function.
@@
In contrast, the four-layer LGMD neural network proposed by Rind et al. appears particularly attractive to robotic engineers. The first implementation of this four-layer LGMD network for collision detection in ground mobile robots occurred in 2000, where a Khepera mobile robot equipped with a CCD camera was used to perceive real-world environments. Collision detection and avoidance were achieved using the simulation software IQR421 [10]. Subsequently, Yue et al. combined genetic algorithms with the LGMD neural network to address collision detection in car driving scenarios [63]. Further improvements were made by incorporating artificial pathway and layer into the four-layer LGMD network, thereby enhancing its performance when managing complex and dynamic environments, alleviating negative impact by camera exposure [62] (See Fig. 5B and Fig. 6A). Building on similar ideas, Yue and Rind continued exploring the potential of LGMD-based near-range path navigation, including the integration of panoramic images and regular CCD images for improved collision detection and avoidance [64, 65]. Moreover, Silva et al. extended the LGMD model by integrating previous work from [62] and [66]. They verified a linear relationship between time-to-collision444time difference between peak firing time of LGMD model (predicted time) and genuine collision time and time-of-collision, then investigated the $l/|v|$ ratio in MATLAB simulations and in online DKRK8000 robot navigation [67].
Another milestone was established in 2017 when the LGMD-based neural network model was built for the first time into the embedded vision system. Hu et al. adapted a similar vision-based LGMD network as in [62] into a small autonomous mobile robot called “Colias[68]. The robot performance was validated through autonomous navigation in an arena mixed with obstacles (see Fig. 6C), demonstrating the reliability and robustness of the embedded vision system of LGMD case. Later, Fu et al. further facilitated collision-free navigation for Colias robot by gradually incorporating spike frequency adaptation, ON and OFF visual pathways, as well as feedback neural computation [13, 69] (See Fig. 5C). In the latest works, the feedback closed-loops in ON/OFF channels are capable of implementing ON/OFF-contrast loom-selectivity of the micro-robot through easily mediating a singular parameter [69, 70]. Except the wheeled robots, a computational study [71] also showcased success of the LGMD1 model as collision detector in a hexapod robot walking through a structured indoor environment (see Fig. 6E).
Regarding aerial robot scenarios, Zhao et al. utilized LGMD-based vision detectors to achieve collision-free flight of a small quadcopter and the model is highly biomimetic that simulates spatiotemporal interaction between excitation and inhibition throughout the pre-synaptic connections (see Fig. 6D)[72, 73, 15]. More recent studies have demonstrated that a variety of LGMD-based vision system also holds great promise for boosting the safe navigation of small UAVs [15, 74]. The UAV mounted with an attention-based LGMD visual system can even detect and flexibly avoid thin power lines during flight [75].
@@
Furthermore, engineers and computational modelers have explored ways to enhance the looming selectivity of LGMD1-based perception methods by incorporating various computational mechanisms. These methods can be broadly classified into two categories: the first approach involves integrating mechanisms such as ON/OFF competition or directional sub-networks to suppress responses against receding or translating stimuli [80, 41]. The second approach employs binocular vision or additional units to estimate changes in the depth of moving objects, thereby distinguishing between approaching, receding, and translating stimuli [66, 81, 82]. However, these mechanisms could essentially increase computational burden, have not yet been validated by real-world robotic implementations.
3.2 LGMD2-based collision detector with enhanced selectivity for ground robot
On the other hand, due to limited biological research on LGMD2 and its selective response to only OFF contrast looming stimuli, i.e. proximity of darker object relative to background, LGMD2 has relatively very limited numerical modeling and applications in robotics. The seminal work on LGMD2 modeling appeared late in 2015 when Fu and Yue introduced ON and OFF visual pathways and non-linearity to simulate the LGMD2’s selectivity for OFF contrast stimuli [31]. According to its specific selectivity that matches well with the visual events faced by ground robots where most foreground moving targets intrinsically become darker than their background, the LGMD2 neural network model was optimized and smoothly embedded into the Colias micro-robot, achieving a collision detection and avoidance success rate of 95.3% in arena tests, showing improved selectivity to approaching over receding and translating stimuli [83].
@@ 3.2 LGMD2-based collision detector with enhanced selectivity for ground robot
As LGMD1 and LGMD2 have different selectivity that could benefit addressing variable real-physical collision challenges, the first hybrid LGMD1-LGMD2 neural model was proposed to implemented as embedded vision system for micro-robot navigation through both dark and bright environments [84]. In this research, the robot’s LGMDs vision behaved similarly to revealed parts of biological neurons that demonstrated the effectiveness of LGMD1 and LGMD2 together for artificial collision-detecting system. Unlike LGMD1, LGMD2 possesses a single large dendritic field, leading to the assumption that FFI is absent in its pre-synaptic circuitry. To accurately simulate the LGMD2’s selectivity, Fu et al. introduced an adaptive inhibition mechanism as a substitute for FFI, thereby aligning with physiological observations and enhancing the performance of the LGMD2-based visual system in real-world tasks including complex car and UAV scenarios [32, 14, 85].
3.3 New application scenarios and sensor techniques with LGMD-based dynamic vision
Researchers are actively seeking opportunities to extend LGMD research to new application scenarios and sensor technologies. A hybrid LGMDs visual system, integrating the firing rate coordination of LGMD1 and LGMD2 neuronal models, was developed as a visual sensing modality for individual micro-robots to enhance intelligent traffic scenarios [86]. This system aims to address automatic collision detection challenges in city and highway traffic simulations. This research represents an initial attempt to evaluate the effectiveness of LGMD-based visual systems in dynamic traffic environments, tractably and at low price, marking a step toward the development of neuromorphic collision sensors for autonomous vehicles. Benefiting from its low computational power, LGMD-based collision detection methods have been combined with path integration mechanisms discovered in sweat bees and desert ants, enabling collision-free autonomous navigation toward specific goal locations, even in the presence of both static and moving obstacles [87, 88, 89].
@@ 3.3 New application scenarios and sensor techniques with LGMD-based dynamic vision
Unlike conventional cameras, event-based cameras detect changes in light intensity asynchronously, producing sparse, high-temporal-resolution data with low power consumption and a wide dynamic range [90]. This makes event-based cameras ideal inputs for motion perception models. When interpreted with spiking neural networks (SNN), event camera-based LGMD models compensate for the limitations of traditional RGB cameras, effectively detecting ultra-fast looming objects or targets in complex and diverse low-light environments, even facilitating the control of UAVs during flights [91, 92, 79]. Recently, a fractional spiking neuron model was proposed to simulate the spiking initiation zone of the LGMD, where ON/OFF events from a dynamic vision sensor are encoded as inputs [92]. This research validates the effectiveness of this sensor strategy in emulating the neuronal dynamics of LGMD. Furthermore, with optimization techniques, Salt et al. utilized an event-driven sensor to develop an LGMD-based UAV obstacle avoidance algorithm, demonstrating that the internal spiking dynamics can be effectively represented within a spiking LGMD neural network [93]. Moreover, thermal cameras have been integrated into LGMD computational models to enhance performance [94]. Unlike previous approaches that relied on normalization mechanisms to boost responses to low-contrast stimuli, Zhang et al. recently introduced thermal imaging as an alternative solution. By leveraging thermal cameras, the ability of LGMD-based models to detect approaching objects was significantly improved, particularly in low-light conditions, offering a candidate method for collision detection in challenging lighting environments.
@@
In summary, LGMD-based collision detection methods provide a distinctive, visually driven, and data-free approach, standing in stark contrast to deep learning and sensor-based techniques. While deep learning models achieve remarkable accuracy by identifying intrinsic patterns within large datasets, they demand significant computational resources, extensive data, additional sensors, and lengthy training processes. In comparison, LGMD-based methods leverage biologically plausible neural structures to enable parsimonious and efficient collision detection without requiring high-performance hardware or large datasets. This makes them particularly well-suited for real-time computing in mobile machines, where timeliness, computational simplicity, and energy efficiency are paramount. Despite these advantages, LGMD-based models face notable challenges in robustness when applied to diverse, variable real-world scenarios. They lack the adaptability and flexibility demonstrated by real LGMD neurons, especially in handling the complexities of dynamic and noisy environments. Addressing these limitations presents significant opportunities for future research to further refine and expand LGMD-based dynamic vision systems and their applications.
@@ 4 How these methods feedback to neuroscience
The performance of LGMD-based robotic systems during complex real-world interaction tasks has directly demonstrated a robustness in looming selectivity comparable to that observed in biological LGMD neurons. In particular, the computational model of LGMD2 has been integrated into the micro-robot - Colias, demonstrating a selective looming response to darker objects during real-world navigation [32]. Additionally, the Strider robot [21], implemented with a four-layer LGMD1 network incorporating a different correlation framework, revealed a linear relationship between time to collision and the size-to-speed ratio ($l/|v|$), closer to previous findings by Gabbiani et al. [36, 37]. F. Claire Rind, a neuroscientist who has dedicated several decades to studying the LGMD neurons, remarked that “Insects, and locust looming detectors in particular, have already provided inspiration for visual control of robots, unmanned autonomous vehicles, and aerial drones.” in a recent discussion on insects’ 3D vision [33]. Her insights underscore the profound impact of bio-inspired research in bridging the gap between neuroscience and technological innovation, demonstrating how understanding simple biological systems can drive advancements in robotics and autonomous systems. Actually, the earlier work based on Rind’s LGMD1 neural network [8], i.e., the Khepera robot, controlled by an LGMD-based collision detection and avoidance system, exhibited reasonable behavior—specifically, as the moving speed increased, the robot generated collision warnings earlier, thereby maintaining a greater distance from obstacles [62].
@@
Bio-inspired robotics have the potential to contribute far beyond mere “verification”. Two decades ago, scholars proposed that bio-robotic studies could serve as effective paradigms for understanding animal behavior, suggesting that the performance of bio-inspired robotics might also offer valuable insights into neuroscience and behavior [95]. Regarding LGMD2, the strongest feedback to neuroscience is the probable existence of ON/OFF-contrast encoding neurons or neural pathways prior to the dendrites of LGMDs [32, 24]. At this point, Rind also commented “Simulations show how response tuning for dark transitions in the LGMD2 could complement LGMD1 responses”[32]. Unlike research into fly visual systems, there is very limited evidence for whether/where the ON/OFF channels exist in locust’s visual circuitry. However, the long-term knowledge of separate channels for ON/OFF-contrast FFI in the LGMD [50], along with recent work on the substantial segregation of ON-excitation to field C of the LGMD1 [16] all informed the significance of polarity motion vision in locust’s visual systems. Accordingly, considering the homology between different insects’ visual brains, and the universal findings of polarity vision across researched animals’ dynamic vision systems, the simulation of LGMD2’s circuits, and separation of ON and OFF based FFI provide strong hypothesis to neuroscience.
@@
In recent studies, the competition between ON/OFF channels also indicate how opposite-polarity signals could interact to suppress translating-induced excitation in order to sharpen up the looming selectivity of LGMD [41]. Furthermore, the performance of LGMD-based real-world applications is partially restricted in noisy and low-contrast environments. Keil’s computational model suggested that threshold-involved neuronal signal integration could pool signals from different channels, thereby enhancing looming selectivity in noisy environments [59]. By analyzing realistic synaptic integration within bio-physically accurate neuronal models, Poleg-Polsky found that introducing an additional threshold for dendritic spikes significantly increased neuronal tolerance to a wide variety of noise [96]. Besides, the recent studies proposed that normalizing photoreceptor signals can enhance LGMD responses to low-contrast stimuli [85, 97]. Similarly, global inhibition observed in the locust’s visual pathways helps maintain excitation within the dynamic range of the neuron, thereby preserving functional selectivity under varying illumination conditions [51]. These mechanisms are also present in other sensory circuits, highlighting their fundamental role in neural processing [98]. Moreover, the deepened LGMD1-LGMD2 cascade network shows selectivity to only approaching targets which might reveal some cooperative mechanisms interacting the pre-synaptic dendrites between LGMD1 and LGMD2, nevertheless unknown in terms of physiology [85, 11].
@@ 5 Discussion
Application demands - First, there is an urgent need for efficient, low-energy collision detection methods in real-world applications. Collisions are ubiquitous and pose significant hazards to both living creatures and intelligent robotic systems. Current collision detection methods are often energy-intensive due to the need for map reconstruction and object recognition [99, 100, 101], or they rely heavily on specific sensors such as radar or wireless ultraviolet light [102, 103]. Insects, with their tiny brains and highly effective collision avoidance behaviors, offer an ideal model for developing parsimonious and efficient collision detection systems. The evolved, specialized structure of looming-selective neural circuits in insects provides a foundation for designing straightforward models capable of addressing real-world collision challenges. There are two identified types of neurons or neuron classes that respond selectively to approaching objects: the LGMDs (one LGMD1/LGMD2 per eye in grasshoppers) and the lobula plate/lobula columnar type-2 (LPLC2) neurons ($\sim 80$ LPLC2 neurons per eye in flies). LPLC2 neurons are located in the lobula complex of Drosophila, and unlike LGMD, their excitation relies on the precise spatial arrangement of dendrites within a directionally selective motion feature map provided by T4/T5 neurons [104, 105]. LPLC2 neurons prefer motion that is outward-from versus inward-toward their receptive field center, and this radial motion opponency causes LPLC2 to be excited by looming objects at the center of the receptive field rather than by receding objects. However, this characteristic limits LPLC2’s response in near-miss scenarios or to looming patterns from different directions, making it potentially insufficient for handling complex real-world challenges. In contrast, LGMD neurons can detect imminent collisions from various directions in both stationary and moving backgrounds, without requiring directionally specific afferent signals [106, 30]. The robustness of LGMD responses makes it a suitable alternative for generating accurate yet efficient collision detection methods.
@@
Biological prominence - Second, comprehensive research in neuroscience and anatomy on the LGMD neuron has significantly contributed to the successful development of artificial systems. The LGMD neuron is large in size, with an extensive dendritic structure that makes it easier to study using various experimental protocols. Dendritic field A spans approximately $7705\mu m$ in length and $108,400\mu m^{2}$ in surface area, while dendritic fields B and C are shorter, with lengths of approximately $4438\mu m$ and $2517\mu m$, surface areas of $28,100\mu m^{2}$ and $13,500\mu m^{2}$, and radii of $0.8\mu m$ and $0.7\mu m$, respectively [107]. These lengths and areas are for the combined length of all dendritic branches. This morphology facilitates research using tools such as extracellular hook electrodes, two-photon microscopy, and optogenetics [18, 107, 108, 109]. Furthermore, investigating the neuronal response of the LGMD is particularly feasible in laboratory settings because approaching objects can be effectively simulated using two-dimensional projections. These factors have sustained the vibrant wave of neuroscience and anatomical research on LGMD neurons since their discovery in the 1970s. Over time, this research has become increasingly comprehensive, addressing aspects ranging from morphology to membrane conductance [110, 111]. Consequently, it has provided an abundant source of inspiration for developing LGMD-based applications in real-world scenarios.
@@
Although the neuronal response of LGMDs can be effectively simulated using the NEURON simulation environment [^5] https://www.neuron.yale.edu/neuron/ [20], the afferent mechanisms driving LGMD responses to different types of motion remain largely unexplored. Olson et al. [22] focused on replicating the neural signal processing of LGMD by incorporating global, lateral, and feedforward inhibition, demonstrating alignment between model outputs and real LMC responses. However, their model was limited in scope, considering only the OFF pathway while excluding the ON pathway, ON/OFF interactions, as well as self-inhibition and lateral excitation. The limited scope covered by Olson et al. is a sign that there is more work to be done towards full processing dynamics of LGMDs.
Moreover, the connectivity between LGMD neurons and other neurons within the locust neuropile remains largely undefined. Recent connectomic data from Drosophila suggested that integrating whole-field looming detection pathways with local directional signals can significantly enhance collision detection accuracy in both simulated and real-world environments [122, 123]. Additionally, modelers have combined LGMD1 with LPLC2 to improve looming perception performance in real-world scenarios, showing significant promise for high-speed collision detection tasks [124, 125]. Another unresolved question concerns the evolutionary rationale behind the emergence of distinct LGMD1 and LGMD2 neurons, each with unique morphology. Understanding why and how these specialized neurons developed over thousands of years is critical for replicating their adaptability and efficiency in engineered models.
@@
While we are still far from fully understanding the intrinsic mechanisms that enable the locust nervous system to achieve its remarkable collision detection capabilities, even a simple replication of the LGMD neuronal response characteristics has already proven sufficient for developing low-energy, robust collision detection methods. This bio-inspired paradigm exemplifies a fundamentally different methodology compared to deep learning-based collision detection strategies. Deep learning methods, though originally inspired by brain functions, primarily emphasize the learning process. They derive their strength from uncovering intrinsic patterns and features within extensive training datasets, achieving outstanding performance through experience-based adaptation. In contrast, LGMD-based collision detection methods focus on directly replicating neuronal responses and neural architectures, providing an alternative approach for real-world collision detection that operates without relying on learning or large datasets. This bio-inspired paradigm of LGMD offers a promising pathway for efficient collision detection, especially in scenarios where simplicity, low computational overhead, and rapid responses are crucial. We believe these two paradigms—deep learning-based and LGMD-based approaches—can complement and enhance each other. Deep learning methods can provide information about approaching objects and the agent’s self-location, further enhancing the performance of LGMD-based collision detection methods, particularly in complex real-world scenarios. Conversely, LGMD-based methods can act as supplementary mechanisms to support deep learning models in situations where training data are limited or unavailable. Together, these approaches form a synergistic framework for advancing collision detection technologies.